# Controllable Graph Generation with Diffusion Models via Inference-Time Tree Search Guidance

### v1

@@ Abstract.
Graph Generation, Diffusion Models, Inference-Time Guidance, Monte Carlo Tree Search
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Graphs are powerful structures for modeling complex relationships among entities (Kipf and Welling, 2017; Hamilton et al., 2017; Veličković et al., 2018; Wang et al., 2024a, 2025c) and form the backbone of numerous Web-related systems, such as knowledge graphs, recommender networks, and online interaction graphs. The ability to generate and manipulate such graphs under domain-specific constraints has become a fundamental problem in graph learning (Bonifati et al., 2020; Simonovsky and Komodakis, 2018; Shi et al., 2020; Vignac et al., 2023a; Liu et al., 2024). For example, in molecular and material design (Liu et al., 2024)—where molecules can be viewed as structured graphs—generation models aim to discover new candidates that satisfy specific chemical or physical properties. Similarly, in broader Web contexts, graph generation underpins applications such as link prediction (Hamilton et al., 2017), knowledge graph completion (Chen et al., 2020), and the simulation of evolving information networks (Myers et al., 2014). Despite these diverse applications, generating graphs that are valid, diverse, and aligned with complex objectives remains a significant algorithmic challenge.
@@ 3. Preliminary
where $\epsilon$ is Gaussian noise and $\epsilon_{\theta}$ is a neural network that predicts the noise component.
In graph generation, diffusion can be applied either directly in the graph space (adjacency and feature matrices) or in a continuous latent space learned by an encoder. Graph-space diffusion preserves structural semantics explicitly, but the discrete and combinatorial nature of graphs makes the denoising process unstable and computationally expensive. Latent-space diffusion alleviates these issues by operating in a continuous space, enabling smoother optimization and more scalable training; however, it often loses fine-grained structural information, leading to limited controllability. This trade-off highlights the need for designs that can better balance structural fidelity and generative flexibility.
@@ 4. Methodology
4.1. Framework
We propose TreeDiff, an MCTS-guided diffusion framework for controllable and stable graph generation. The core idea is to view the denoising process as a sequential decision problem and augment it with structured search, so that alternative denoising trajectories can be explored, compared, and optimized during inference. While MCTS provides a natural mechanism for long-term planning, directly applying it to diffusion sampling is infeasible due to three key obstacles: (i) the enormous depth of diffusion trajectories, (ii) the tension between efficiency in latent space and stability in graph space, and (iii) the prohibitive cost of evaluating complete rollouts.
To address these issues, TreeDiff incorporates three complementary components (Figure 3). First, a macro-step expansion strategy reduces tree depth by grouping multiple denoising steps into a single node transition. Second, a dual-space denoising mechanism couples latent-space denoising with graph-space guidance, balancing efficiency with structural fidelity. Third, a dual-space verifier combines fast validity checks with more expensive property-based objectives, enabling efficient yet reliable reward estimation. These designs allow TreeDiff to allocate inference-time computation adaptively, achieving improvements in graph quality and controllability. The implementation details are provided in Appendix C.
@@ 4.2. Macro-Step Expansion
In MCTS, the Expansion phase controls how new nodes are added to the search tree. For diffusion-based graph generation, each reverse step only makes a minor refinement to the current graph state, yielding an extremely deep, fine-grained denoising trajectory. If expansion were performed at every diffusion step, the tree would contain hundreds or thousands of layers with near-duplicate parent/child states, making exploration prohibitively slow and preventing meaningful long-horizon planning. It is therefore crucial to redefine the granularity of expansion so that the tree grows over coarser, semantically meaningful transitions rather than raw step-level perturbations. We introduce a macro-step expansion strategy that groups multiple reverse denoising updates into a single transition. Instead of adding a child for each step $t\!\to\!t-1$, TreeDiff expands from ${\mathcal{G}}_{t}$ directly to ${\mathcal{G}}_{t-k}$ for a chosen horizon $k$, thereby preserving the overall denoising dynamics while drastically reducing the effective search depth. From the MCTS perspective, this turns expansion from fine-grained local moves into a hierarchical exploration of the denoising landscape.
Formally, let ${\mathcal{G}}_{t}$ denote the intermediate graph at reverse step $t$. A single reverse update, induced by the learned noise predictor $\epsilon_{\theta}({\mathcal{G}}_{t},t)$ and variance schedule $\{\beta_{t}\}_{t=1}^{T}$, can be written as a sampler ${\mathcal{G}}_{t-1}\;=\;\Phi_{\theta}({\mathcal{G}}_{t},t;\,\xi_{t})$, where $\Phi_{\theta}$ denotes the one-step reverse transition and $\xi_{t}\sim\mathcal{N}(0,{\mathbf{I}})$ is the step-wise randomness. We define a $k$-step macro-step operator by composing $k$ such updates:
${\mathcal{G}}_{t-k}$ $=\Phi_{\theta}^{(k)}\big({\mathcal{G}}_{t},t;\xi_{t-k+1:t}\big)$
@@
with $\xi_{t-k+1:t}=\{\xi_{t-k+1},\ldots,\xi_{t}\}$. During Expansion, TreeDiff samples a small set of candidate macro-steps and inserts the resulting as children of ${\mathcal{G}}_{t}$ in the search tree $\mathcal{T}$}.
Adaptive Step Sampling. To further enhance exploration diversity, TreeDiff dynamically determines the macro-step length $k$ at each expansion rather than fixing it across the tree. For a node corresponding to step $t$, we estimate the remaining denoising steps $T_{\text{rem}}=t$ and the remaining allowable tree depth $D_{\text{rem}}$. A base step size is computed as $k_{\text{base}}=\frac{T_{\text{rem}}}{D_{\text{rem}}}$, representing the average number of diffusion updates each future expansion should roughly cover. Then, for every expansion, TreeDiff samples the actual $k$ from a Gaussian distribution centered at this estimate:
(4) $k\sim\mathcal{N}\!\big(k_{\text{base}},\,(\sigma_{k}\,k_{\text{base}})^{2}\big),$
and clips it to ensure $1\leq k\leq T_{\text{rem}}$, where $\sigma_{k}$ is a deviation coefficient (set to $0.1$ by default). This stochastic resampling introduces controlled randomness in temporal granularity: some branches advance aggressively with longer macro-steps, while others refine locally with shorter ones. When the remaining steps fall below $k_{\text{base}}$, TreeDiff performs the remaining denoising to reach ${\mathcal{G}}_{0}$. This adaptive sampling further balances exploration and exploitation in MCTS, leading to more robust and diverse denoising trajectories.
@@ 4.3. Dual-Space Denoising
While macro-step expansion enables TreeDiff to explore diffusion trajectories efficiently, it leaves open a key question for MCTS: in which representation should Selection and Expansion operate? Each node in the search tree corresponds to a partially denoised graph, which can be expressed either in the continuous latent space or in the discrete graph space. Choosing either space alone introduces a fundamental trade-off. Operating purely in the latent space provides smooth, differentiable transitions suitable for selection policies such as UCB (Auer et al., 2002) or PUCT (Silver et al., 2016), but decoded trajectories often drift into invalid graphs, undermining reliable value estimation. Conversely, expanding and evaluating directly in the discrete graph space preserves structural validity and enables domain-specific scoring (e.g., chemical properties), but leads to combinatorial branching and discontinuous transitions, making selection noisy and unstable. This mismatch between continuous exploration and discrete evaluation disrupts MCTS’s ability to balance exploitation and exploration across the denoising trajectory. To bridge this gap, TreeDiff introduces a dual-space denoising mechanism that couples continuous latent diffusion with lightweight discrete correction in graph space.
Specifically, dual-space denoising governs the internal dynamics of each MCTS expansion with $k$ steps by weaving continuous diffusion and discrete refinement into a coherent update rule. Let ${\mathcal{G}}_{t}^{\text{latent}}$ denote the latent representation of a partially denoised graph at reverse step $t$, and ${\mathcal{G}}_{t}^{\text{structure}}$ its decoded discrete counterpart. TreeDiff first advances $n$ latent denoising steps to reach ${\mathcal{G}}_{t-n}^{\text{latent}}$, then projects this intermediate state into the structural domain, ${\mathcal{G}}_{t-n}^{\text{structure}}=\mathrm{Dec}({\mathcal{G}}_{t-n}^{\text{latent}},t-n)$, where a lightweight graph denoiser $p_{\psi}$ performs $m$ categorical refinement steps ($m\ll n$) to enforce topological consistency and inject symbolic priors. The refined graph ${\mathcal{G}}_{t-n-m}^{\text{structure}}$ is re-encoded into a structure-aware latent ${\mathcal{G}}_{t-n-m}^{\text{latent}}=\mathrm{Enc}({\mathcal{G}}_{t-n-m}^{\text{structure}},t-n-m)$, embedding discrete priors back into the continuous manifold.
@@
4.4. Dual-Space Verifier
In MCTS, the Simulation and Backpropagation phases estimate and propagate the expected reward of each trajectory. Fully rolling out every branch to obtain final graph properties is prohibitively expensive when diffusion involves hundreds of steps and multiple objectives. To address this, TreeDiff introduces a learned dual-space verifier that predicts long-term rewards from partially denoised graphs, replacing costly explicit rollouts. The verifier $V_{\phi}$ is trained on a small set of full trajectories, learning to approximate the terminal value $R({\mathcal{G}}_{0})$ directly from mid-trajectory states. At each reverse step $t$, it takes both the latent representation ${\mathcal{G}}_{t}^{\text{latent}}$ and the decoded discrete structure ${\mathcal{G}}_{t}^{\text{structure}}$ as input and outputs $\hat{R}_{t}\approx R({\mathcal{G}}_{0})$. By jointly reasoning over latent and graph spaces, $V_{\phi}$ remains consistent with dual-space denoising and robust to decoding noise. During search, $\hat{R}_{t}$ replaces full rollouts for scoring and backpropagation, enabling efficient planning under multi-objective settings.
@@ 4.4. Dual-Space Verifier
Tradition
Graph-GA (Jensen, 2019) 85.85 98.05 0.963 46.90 89.50 95.09 1.208 30.15 89.93 96.61 0.984 60.35 7.3
@@
Diffusion
DiGress (Vignac et al., 2023a) 88.62 69.42 2.068 50.61 90.98 68.05 2.366 65.36 91.94 85.62 1.922 53.35 9.1
@@
Inference-Time
Best-of-N (Ma et al., 2025) 87.29 88.50 0.480 91.01 89.29 92.41 0.420 93.80 88.54 95.61 0.360 97.01 5.8
@@
Inference-Time
Best-of-N(Ma et al., 2025) 16.23 109.97 58.11 67.08 6.27 13.54 31.07 359.91 110.13 175.11 14.59 24.07 4.2
@@
The dual-space verifier $V_{\phi}$ is trained as a regression model that predicts the terminal reward $R({\mathcal{G}}_{0})$ from an intermediate state $({\mathcal{G}}_{t}^{\text{latent}},{\mathcal{G}}_{t}^{\text{structure}})$. Training data are obtained from a small set of full denoising rollouts $\{{\mathcal{G}}_{t}\}_{t=0}^{T}$ generated by a pretrained diffusion model, each annotated with the corresponding final reward $R({\mathcal{G}}_{0})=\sum_{k}w_{k}f_{k}({\mathcal{G}}_{0})$ that combines multiple property scores. To improve data diversity and robustness, we further apply a stochastic augmentation procedure: for each latent graph ${\mathcal{G}}_{t}^{\text{latent}}$, we inject Gaussian perturbations $\tilde{{\mathcal{G}}}_{t}^{\text{latent}}={\mathcal{G}}_{t}^{\text{latent}}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma_{a}^{2}{\mathbf{I}})$, and decode $\tilde{{\mathcal{G}}}_{t}^{\text{latent}}$ into $\tilde{{\mathcal{G}}}_{t}^{\text{structure}}$ using $\mathrm{Dec}(\cdot)$. These perturbed pairs $(\tilde{{\mathcal{G}}}_{t}^{\text{latent}},\tilde{{\mathcal{G}}}_{t}^{\text{structure}})$ are included alongside the original rollouts to enrich supervision across local neighborhoods of the latent manifold. The verifier minimizes a mean-squared error objective:
@@
In practice, $V_{\phi}$ adopts a dual-branch architecture: a graph transformer encodes ${\mathcal{G}}_{t}^{\text{structure}}$, a lightweight MLP processes ${\mathcal{G}}_{t}^{\text{latent}}$, and a cross-attention layer fuses the two embeddings before a final projection to the scalar estimate $\hat{R}_{t}$. This design allows the verifier to aggregate global structural patterns and latent diffusion context jointly. During MCTS search, $\hat{R}_{t}$ serves as the node value used for both Simulation and Backpropagation, enabling efficient evaluation of candidate trajectories without full rollouts.
The dual-space verifier closes the search loop in TreeDiff by providing a principled, efficient substitute for full rollouts. Together with macro-step expansion and dual-space denoising, it allows MCTS to operate entirely in latent space while remaining grounded in discrete graph semantics. Moreover, since $R({\mathcal{G}}_{0})$ can represent an arbitrary weighted combination of domain-specific objectives, $V_{\phi}$ naturally supports controllable and multi-objective generation without retraining. This design endows TreeDiff with flexible inference-time reasoning, effectively transforming diffusion sampling into a structured search process guided by learned value estimation.
@@ 5.1. Conditional Graph Generation
Trad.
MoFlow (Zang and Wang, 2020) 91.36 4.467 0.0169 0.145 63.11 20.931 0.0455 0.013 10.4
@@
Inference-Time
Best-of-N (Ma et al., 2025) 99.22 0.153 0.0003 0.912 98.84 3.803 0.0013 0.485 4.8
@@
Trad.
ENF (Garcia Satorras et al., 2021) 39.4 85.0 4.9 16.7
@@
Inference-Time
Best-of-N (Ma et al., 2025) 93.9 98.6 87.5 85.4 4.0

### v2

@@ Abstract.
Graph Generation, Diffusion Models, Monte Carlo Tree Search
copyright: acmlicensedjournalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792489isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Data miningccs: Applied computing Computational biologyccs: Computing methodologies Machine learning approaches
@@ 1. Introduction
Graphs are powerful structures for modeling complex relationships among entities (Kipf and Welling, 2017; Hamilton et al., 2017; velicković2018graph; Wang et al., 2024a, 2025c) and form the backbone of numerous Web-related systems (Zhao et al., 2021; Ju et al., 2022), such as knowledge graphs, recommender networks, and online interaction graphs. The ability to generate and manipulate such graphs under domain-specific constraints has become a fundamental problem in graph learning (Bonifati et al., 2020; Simonovsky and Komodakis, 2018; Shi et al., 2020; Vignac et al., 2023a; Liu et al., 2024). For example, in molecular and material design (Liu et al., 2024)—where molecules can be viewed as structured graphs—generation models aim to discover new candidates that satisfy specific chemical or physical properties. Similarly, in broader Web contexts, graph generation underpins applications such as link prediction (Hamilton et al., 2017), knowledge graph completion (Chen et al., 2020), and the simulation of evolving information networks (Myers et al., 2014). Despite these diverse applications, generating graphs that are valid, diverse, and aligned with complex objectives remains a significant algorithmic challenge.
@@ 3. Preliminary
where $\epsilon$ is Gaussian noise and $\epsilon_{\theta}$ is a neural network that predicts the noise component. In graph generation, diffusion can be applied either directly in the graph space (adjacency and feature matrices) or in a continuous latent space learned by an encoder.
@@ 4. Methodology
We propose TreeDiff, an MCTS-guided diffusion framework for controllable and stable graph generation. TreeDiff formulates diffusion sampling as a sequential decision process and applies structured tree search to optimize denoising trajectories at inference time. Directly combining MCTS with diffusion is challenging due to three factors: (i) extremely long denoising horizons, (ii) the efficiency–stability trade-off between latent and graph spaces, and (iii) the high cost of full-trajectory evaluation. To overcome these challenges, TreeDiff introduces three components (Figure 3): (1) macro-step expansion to compress long trajectories into coarse transitions, (2) dual-space denoising to couple efficient latent updates with graph-level structural correction, and (3) a dual-space verifier to estimate long-term rewards without full rollouts. These designs enable adaptive inference-time computation and substantially improve generation controllability. Implementation details are provided in Appendix C.
@@ 4.1. Macro-Step Expansion
In MCTS, the Expansion phase determines how new nodes are added to the search tree. For diffusion-based generation, each reverse step only slightly refines the graph state, resulting in extremely deep trajectories. Expanding at every step would yield hundreds or thousands of near-duplicate layers, making exploration inefficient and long-horizon planning impractical. We redefine the expansion granularity using a macro-step expansion strategy that groups multiple denoising steps into a single transition. Instead of expanding from ${\mathcal{G}}_{t}$ to ${\mathcal{G}}_{t-1}$, TreeDiff directly expands to ${\mathcal{G}}_{t-k}$ for a horizon $k$, substantially reducing tree depth while preserving overall denoising dynamics. This converts expansion from step-wise local moves into hierarchical exploration over the denoising trajectory.
Formally, let ${\mathcal{G}}_{t}$ denote the intermediate graph at reverse step $t$. A single update is written as ${\mathcal{G}}_{t-1}=\Phi_{\theta}({\mathcal{G}}_{t},t;\xi_{t})$, where $\Phi_{\theta}$ is the learned reverse transition and $\xi_{t}\!\sim\!\mathcal{N}(0,{\mathbf{I}})$. A $k$-step macro transition is defined by composition:
${\mathcal{G}}_{t-k}$ $=\Phi_{\theta}^{(k)}({\mathcal{G}}_{t},t;\xi_{t-k+1:t})$
@@
with $\xi_{t-k+1:t}=\{\xi_{t-k+1},\ldots,\xi_{t}\}$. During Expansion, TreeDiff samples a small set of candidate macro-steps and inserts the resulting states as children of ${\mathcal{G}}_{t}$ in the search tree $\mathcal{T}$}.
Adaptive Step Sampling. Rather than fixing $k$, TreeDiff adapts the adaptive expansion. Given remaining diffusion steps $T_{\text{rem}}=t$ and remaining tree depth $D_{\text{rem}}$, we set $k_{\text{base}}=\frac{T_{\text{rem}}}{D_{\text{rem}}}$, and sample
(4) $k\sim\mathcal{N}\!\big(k_{\text{base}},(\sigma_{k}k_{\text{base}})^{2}\big),\quad 1\leq k\leq T_{\text{rem}},$
where $\sigma_{k}=0.1$. This stochastic scheduling enables both coarse long-range moves and fine local refinement, improving the exploration–exploitation trade-off. When $T_{\text{rem}}<k_{\text{base}}$, the process directly denoises to ${\mathcal{G}}_{0}$.
@@ 4.2. Dual-Space Denoising
Macro-step expansion enables efficient exploration but raises a key question for MCTS: which representation should be used for selection and expansion? Each node can be represented in either latent or graph space, yet both have limitations. Latent-space search provides smooth transitions for UCB (Auer et al., 2002) or PUCT (Silver et al., 2016), but decoded trajectories may become invalid. Graph-space search preserves structural validity and enables property evaluation, but introduces combinatorial branching and discontinuous transitions, leading to unstable selection. This mismatch hinders effective long-horizon planning. To address it, TreeDiff adopts dual-space denoising, coupling continuous latent diffusion with lightweight graph-space correction.
Specifically, dual-space denoising governs each $k$-step expansion by alternating continuous diffusion and discrete refinement. Let ${\mathcal{G}}_{t}^{\text{latent}}$ and ${\mathcal{G}}_{t}^{\text{structure}}$ denote the latent and decoded graph states at step $t$. TreeDiff first performs $n$ latent denoising steps to obtain ${\mathcal{G}}_{t-n}^{\text{latent}}$, then decodes it as ${\mathcal{G}}_{t-n}^{\text{structure}}=\mathrm{Dec}({\mathcal{G}}_{t-n}^{\text{latent}},t-n)$. A lightweight denoiser $p_{\psi}$ applies $m\!\ll\!n$ categorical refinement steps to enforce topological consistency, producing ${\mathcal{G}}_{t-n-m}^{\text{structure}}$, which is re-encoded as ${\mathcal{G}}_{t-n-m}^{\text{latent}}=\mathrm{Enc}({\mathcal{G}}_{t-n-m}^{\text{structure}},t-n-m)$.
@@
Tradition Graph-GA (Jensen, 2019) 85.85 98.05 0.963 46.90 89.50 95.09 1.208 30.15 89.93 96.61 0.984 60.35 7.3
@@
Diffusion DiGress (Vignac et al., 2023a) 88.62 69.42 2.068 50.61 90.98 68.05 2.366 65.36 91.94 85.62 1.922 53.35 9.1
@@
Inference-Time Best-of-N (Ma et al., 2025) 87.29 88.50 0.480 91.01 89.29 92.41 0.420 93.80 88.54 95.61 0.360 97.01 5.8
@@
Inference-Time Best-of-N(Ma et al., 2025) 16.23 109.97 58.11 67.08 6.27 13.54 31.07 359.91 110.13 175.11 14.59 24.07 4.2
@@
4.3. Dual-Space Verifier
In MCTS, the Simulation and Backpropagation phases estimate long-term rewards for candidate trajectories. However, full rollouts are prohibitively expensive for diffusion with hundreds of steps and multiple objectives. TreeDiff therefore introduces a learned dual-space verifier $V_{\phi}$ to predict terminal rewards from partially denoised states. At step $t$, it takes both the latent representation ${\mathcal{G}}_{t}^{\text{latent}}$ and the decoded structure ${\mathcal{G}}_{t}^{\text{structure}}$ as input and outputs $\hat{R}_{t}\approx R({\mathcal{G}}_{0})$. This replaces explicit rollouts during search, enabling efficient multi-objective planning while remaining consistent with dual-space denoising.
The verifier is trained as a regression model on full denoising trajectories $\{{\mathcal{G}}_{t}\}_{t=0}^{T}$ generated by a pretrained diffusion model, each labeled with the terminal reward $R({\mathcal{G}}_{0})=\sum_{k}w_{k}f_{k}({\mathcal{G}}_{0})$. To improve robustness, we apply stochastic augmentation by perturbing latent states $\tilde{{\mathcal{G}}}_{t}^{\text{latent}}={\mathcal{G}}_{t}^{\text{latent}}+\epsilon$, $\epsilon\sim\mathcal{N}(0,\sigma_{a}^{2}{\mathbf{I}})$, and decoding them into $\tilde{{\mathcal{G}}}_{t}^{\text{structure}}$. The verifier minimizes
@@ 4.3. Dual-Space Verifier
In practice, $V_{\phi}$ adopts a dual-branch architecture: a graph transformer encodes ${\mathcal{G}}_{t}^{\text{structure}}$, an MLP processes ${\mathcal{G}}_{t}^{\text{latent}}$, and cross-attention fuses both before predicting $\hat{R}_{t}$. During MCTS, $\hat{R}_{t}$ is used as the node value for simulation and backpropagation, enabling efficient evaluation without rollouts.
The dual-space verifier completes TreeDiff’s search loop by providing fast, reliable value estimates. Together with macro-step expansion and dual-space denoising, it enables latent-space planning while remaining grounded in discrete graph semantics. Since $R({\mathcal{G}}_{0})$ can combine arbitrary objectives, the verifier naturally supports controllable multi-objective generation at inference time.
@@ 5.1. Conditional Graph Generation
Trad. MoFlow (Zang and Wang, 2020) 91.36 4.467 0.0169 0.145 63.11 20.931 0.0455 0.013 10.4
@@
Inference-Time Best-of-(Ma et al., 2025) 99.22 0.153 0.0003 0.912 98.84 3.803 0.0013 0.485 4.8
@@
Trad. ENF (Garcia Satorras et al., 2021) 39.4 85.0 4.9 16.7
@@
Inference-Time Best-of-N (Ma et al., 2025) 93.9 98.6 87.5 85.4 4.0
@@ 6. Conclusion
Acknowledgements.
The work of Z. Wang, C. Zhang, and Y. Ye was partially supported by the NSF under grants IIS-2533550, IIS-2321504, IIS-2217239, CNS-2426514, and CMMI-2146076, Notre Dame Strategic Framework Research Grant (2025), and Notre Dame Poverty Research Package (2025). Any expressed opinions, findings, and conclusions or recommendations are those of the authors and do not necessarily reflect the views of the sponsors.